List all subsets of the set {8, 1, 3}
List the elements of the given set in roster notation. {x l7 - x=9; an integer}

Answer by MathLover1(20850)


there is 8 subsets
{} {8} {1} {3} {8,1} {8,3} {1,3} {8,1,3}

List the elements of the given set in roster notation. {x |7 - x=9; an integer}
7+-+x=9
7+-+9=x
-2=x
the elements of the given set: {-2}